OpenStudy (anonymous):

f(x)=3ln(x) Find inverse function & the domain of inverse.

OpenStudy (anonymous):

to find inverse: let f(x) = y and invert x and y. so y = 3ln(x) becomes x = 3ln(y). Now solve for y to put it back in the form of a f(x) ln(y) = x/3 inverse operation of ln(y) is e^y e^(ln(y)) = e^(x/3) y = e^(x/3) = g(x) = inverse of f(x)

OpenStudy (anonymous):

oh okay thank you!

OpenStudy (anonymous):

inverse operation of ln(y) is e^y e^(ln(y)) = e^(x/3)

OpenStudy (anonymous):

I ment this step

OpenStudy (anonymous):

it is just the definition of the logarithmic function. For: \[y = \log_b(x)\] \[x = b^y\] \[\ln(y) = \log_e(y)\]

OpenStudy (anonymous):

Wait but right now my paper says x/3=e^y right

OpenStudy (anonymous):

So would it be log e x/3=y

OpenStudy (anonymous):

that is the truth. those two equations are equal

OpenStudy (anonymous):

oh. i think i see the misunderstanding: the point of that step was that: \[\ln(e^y) = y\]

OpenStudy (anonymous):

which is what i meant by inverse operation same way that x = sin^(-1)(y) is the inverse of y = sin(x), used to find the angle x

OpenStudy (anonymous):

and that sqrt is inverse of square

OpenStudy (anonymous):

So would it be

OpenStudy (anonymous):

log e x/2 = y?

OpenStudy (anonymous):

yup. assuming that 2 is a typo :)